Let’s suppose you (USA dealer) imported a product from German on Dec 1, 2018 at € 300, payable in 60 days. You sold all of them in the US market at $400 in cash on Dec 15, 2018. The company's fiscal year ends on Dec 31. You paid to your German supplier on Feb 1, 2019. Below, please find the exchange rate information: Dec 1, 2018: 1.2 $/€. Dec 31, 2018: 1.5 $/€ Feb 1, 2019: 1.0 $/€ What was net income for 2018 and 2019, respectively?

Group of answer choices

A) -$50, $150

B) $250, -100

C)$200, $100

D)$250, $100
Answer #1

2018
Purchase Cost = E300 x $.12 = $360
Sales = $400
Exchange loss = (1.5-1.2) x 300 = $90

Net Income = Sales - Purchase cost - Exchange loss
= $400 - $360 - $90 = ($50)

2019
Exchange gain = (1.5-1) x 300 = $150
Net Income = $150

Answer is A) -$50, $150